OpenStudy (anonymous):

how do i simplify (2d^2 +14d +24) / (d^2 -9)

OpenStudy (anonymous):

factorize the denominator and numerator

OpenStudy (anonymous):

then cancel out the common terms

OpenStudy (anonymous):

how do i factorize the denominator and numerator?

OpenStudy (anonymous):

d^2 - 9 = (d+3)(d-3)

OpenStudy (anonymous):

and 2d^2 + 14d + 24 = 2(d+4)(d+3)

OpenStudy (anonymous):

now d+3 is common to both so it cancels out...so it comes to 2(d+4)/(d-3)

OpenStudy (anonymous):

can you help me with a couple more?

OpenStudy (anonymous):

go on...try the others on ur own lyk i told u 2